The length of a rectangle is 6 cm less than twice it's width. Find the dimesions of the rectangle if the area is 108cm^2
Answer by checkley77(12844)

L=2W-6
LW=108
(2W-6)W=108
2W^2-6W-108=0
2(W^2-3W-54)=0
2(W-9)(W+6)=0
W-9=0
W=9 ANS. FOR THE WIDTH.
L=2*9=6
L=18-6
L=12 ANS. FOR THE LENGTH.
PROOF:
9*12=108
108=108